OpenStudy (anonymous):

please help! The midpoint of a line segment with end points as (-1, 5) and (x2, 7) is (-4, 6). What is the value of x2?

OpenStudy (anonymous):

((-1,5) + (x2, 7))/2 = (-1+x2, 12)/2 = (-4, 6) so -1 + x2 = -4 *2 and then x2 = -7

OpenStudy (anonymous):

thank you very much!! (:

OpenStudy (anonymous):

welcome:)